OpenStudy (anonymous):

determine the center and the radius of the circle: x^2+y^2-24x-16y-17=0

OpenStudy (anonymous):

solve by completing the square twice x^2-24x+144+y^2-16y+64 = 144+64+17

OpenStudy (anonymous):

(x-12)^2+(y-8)^2 = 225

OpenStudy (anonymous):

center (12, 8) radius Sqrt[225]=15